# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ Abstract
We derive the energy-differential cross section and energy loss rate for dissipative self-interacting dark matter (dSIDM) models within the Born regime using perturbative quantum field theory. Six dissipative scenarios are considered, incorporating the emission of particles that may be either massless or possess a kinematically allowed light mass. Both short-range and long-range force-mediated dSIDM interactions are examined. In the non-relativistic regime we obtain closed-form expressions of the energy-differential cross sections by a controlled expansion in the initial relative dark matter velocity. Up to trivial factors, the leading-order squared emission amplitude is model-independent for massless emissions. Model dependence arises for massive particle emission and at the next-to-leading order. The latter reduces to three distinct cases. The derived analytical expressions for dipole and quadrupole emission exhibit excellent agreement with numerical computations, providing simple, ready-to-use formulas. These expressions are benchmarked against full numerical results to confirm their accuracy. Furthermore, we analyze the behavior of these processes in the soft emission limit. Our results show that additional corrections are necessary when applying factorization to ensure consistency between the soft energy-differential cross section and the full counterparts across a broad energy range. Finally, we investigate the regime of perturbative validity in terms of the model parameters, identifying the conditions under which our results are applicable.
@@ 1 Introduction
Relaxing the assumption of a collisionless DM component also opens the door to additional phenomena, particularly dissipation, additionally modifying the thermodynamics of halos: radiative channels shorten the cooling time relative to dynamical and relaxation times, enabling contraction and catalyzing collapse to black holes [31, 32, 33], compact substructure [34, 35, 36], or the formation of rotationally supported dark disks [37, 38, 39, 40]. Because of constraints on the self-interaction cross sections $\sigma/m$, typically only a fraction of the total DM is assumed to participate in such dissipative sector. Among other models, a well-studied dissipative realization is Atomic Dark Matter (ADM), which mimics the SM baryonic sector by allowing for the formation of hydrogen-like bound states [41, 42, 43, 44, 45]. Although ADM offers a rich phenomenological playground, much of its astrophysical impact stems from the net energy loss within halos, rather than from the specific details of the atomic spectrum. Consequently, some minimal dissipative models have been put forward in which the DM is simply composed of one candidate (fermionic or scalar) associated with either a massless, unbroken dark photon $U(1)_{D}$ [46, 47], or through a massive dark gauge boson [48, 49]. These stripped-down setups already offer valuable insight into small-scale structure formation, avoiding the large parameter space inherent to more complex models. For additional works on dissipative SIDM (dSIDM) see [50, 51, 52, 53, 54] and references therein.
@@
In the SM, the Bremsstrahlung process corresponds to the emission of a photon in the scattering of two electrically charged particles, and is one of the most fundamental radiative processes. Two distinct cases arise. First, when the particles have different charge-to-mass ratios, dipole emission occurs. This scenario was first treated classically by Kramers [55], followed by quantum mechanical (Born) calculations by Oppenheimer [56] and Sugiura [57], and later by Sommerfeld who derived the exact non-relativistic solution using confluent hypergeometric functions [^1] To avoid the cumbersome exact solution, Elwert introduced a multiplicative correction that extends the Born approximation [58]. [59], a treatment now standard in advanced textbooks. More recently, dipole emissions have been revisited in [60]. Second, when the particles share the same mass-to-charge ratio, leading to the vanishing of the dipole piece and the dominance of quadrupole radiation. The corresponding cross-section was first derived in a Born calculation by Gould [61, 62], and revisited recently in a calculation exact at all orders in the Coulomb interaction of two non-relativistic particles in [63]. Beyond the full energy spectrum, the soft-emission limit is also of particular interest: in this regime, for photon and graviton, the amplitude factorizes into a universal soft factor and a process-dependent hard amplitude [64].
@@
Bremsstrahlung-like processes are also found in many astrophysical and beyond the SM contexts. In stellar environments, similar types of radiation can arise from axion emission [65, 66, 67], neutrino pair production [68, 69, 70], or scalar emission [71, 72]. Unsurprisingly, these multipole emissions also reappear in dissipative dark-sector settings: dipole bremsstrahlung governs cooling in ionized dark-electron–dark-proton plasmas, while quadrupole emission dominates in dark-electron–dark-electron interactions. However, a comprehensive theoretical framework for Bremsstrahlung in generic dark sectors, departing from a simple gauge boson emission that includes dipole and higher multipoles emission, distinguishable and identical scatterers, different emitted quanta, and soft emission, has yet to be fully established.
@@
Motivated by this gap and the diversity of the SIDM proposals, we establish in this work a set of representative minimal dissipative scenarios dominated by Bremsstrahlung emission. Our DM candidates include fermions (Dirac or Majorana) and scalars (real or complex), coupled through either scalar or vector mediator. Two distinct species are included in each scenario, allowing the study of both distinguishable and identical scatterers, and both the long-range (light mediator) and short-range (heavy mediator) regimes are covered. We work throughout in the non-relativistic (NR) limit, appropriate for small-scale structures where the characteristic DM relative velocity is $v_{i}\ll 1$. Within this framework, analytic formulae are derived for the energy-differential cross-section and corresponding energy-loss rate in every scenario. The results are compiled into a single catalogue and validated against numerical calculations and existing literature. Where previous studies relied on process-specific approximations, this catalogue offers a single lookup table covering the parameter space for the most relevant SIDM scenarios, and ready-to-use expressions that can be readily adapted to any combination of couplings, masses, and mediator properties. The present study is restricted to the perturbative Born regime; a detailed investigation of non-perturbative effects will be addressed in a future work.
@@
The paper is organized as follows: In Sec. 2, we introduce the set of dissipative self-interacting dark matter (dSIDM) scenarios we consider in this work, summarize the main existing constraints on the corresponding SIDM parameters, and discuss the domains of validity of our results. In Sec. 3, treating separately long-range and short-range mediation, we first identify the dissipative channels that appear in the Born (perturbative) regime and analyse their relevance in the NR limit, making explicit the emergence of a hierarchy of scales among them. Next, we provide a concise textbook review of NR radiation, distinguishing dipole from quadrupole emission. The section concludes with explicit expressions for the squared emission amplitudes and the energy-differential cross-sections, both for distinguishable and identical particle scattering. In Sec. 4, we investigate the low-energy behavior of the different emission amplitudes and energy differential cross-sections obtained. We employ the standard factorisation that expresses the former as an elastic part multiplied by an emission piece, and discuss the limitations of such a method. We conclude this paper with Sec. 5, where we derive energy loss rates. The Appendices provide further details on our calculations, including the case where the emitted particle carries a small, kinematically allowed mass.
@@ 2.1 Field content and considered interactions
Again, we have a dimensionful trilinear interaction constant $A_{\phi}$ for the real scalar mediator and self-interation quartic couplings $\lambda_{\phi}$ and $\lambda_{S}$. We take all couplings as real. [^2] We exclude off-diagonal couplings such as $\lambda_{12}S_{1}S_{2}\phi$, inducing a decay channel $S_{1}\to S_{2}\phi$ (or $S_{2}\to S_{1}\phi$ when $m_{1}>m_{2}+m_{\phi}$), altering the relative DM populations and introducing tangential complications.
A comment is in order about the Lagrangians (2.3), (2.4), and (2.5) when considering indistinguishable particle scattering, for which we then drop any generational index $j$ on the fields and couplings. In this case, we identify $S_{1}=S_{2}=S$ (and similarly for other fields) and it is understood that the sums over $j$ in (2.3) and (2.4) should be dropped. In other words, the normalization of coupling constants remains unchanged and do not receive a factor of two. Finally, the quartic scalar DM interaction term in (2.5) becomes $(1/4)\lambda_{S}|S|^{4}$ for $S$ and $(1/4!)\tilde{\lambda}_{S}|\tilde{S}|^{4}$ for $\tilde{S}$. It is for this reason that we have not included additional quartic $|S_{j}|^{4}$ interactions in (2.5).
@@
Of course, the possibilities considered above are not exhaustive, but still represent a fair fraction of considered dSIDM possibilities. For simplicity, we shall also only consider exclusive cases where DM is of one principal fermionic or bosonic type. Mixed cases are, however, closely related since, as we shall see, the currents between fermions and scalars share many similarities in the NR limit. Notably absent from the above Lagrangians are axial-vector couplings and pseudoscalar couplings; in the latter case, the Bremsstrahlung emission has been extensively studied in the context of axion emission from stars [65, 66, 67].
With the introduced field content, we identify six principal dSIDM combinations. For the processes where the interaction is mediated by a vector boson and involves the emission of a boson of the same type, two types of DM candidates are relevant for the corresponding dissipative processes: $\chi$ and $S$. These give rise to the dSIDM scenarios $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$, and $S_{1}S_{2}\to S_{1}S_{2}V$. In contrast, when the interaction is mediated by a scalar boson, and is associated with the emission of a scalar particle, the four DM candidates $S_{j}$$\tilde{S}_{j}$$\chi_{j}$, or $\tilde{\chi}_{j}$ are considered. These lead to the dSIDM scenarios $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$,  $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$ and $S_{1}S_{2}\to S_{1}S_{2}\phi$,  $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$.
@@ 2.2 Self-scattering constraints on model parameters
The success of SIDM lies in its ability to remain consistent across a wide range of astrophysical systems. In particular, the associated cross-section per unit of DM mass ${\sigma}/{m_{\chi}}$ can be large enough to alter the inner density profiles of dwarf galaxies, yet small enough to avoid constraints on galaxy and galaxy cluster scales. Consequently, the principal bounds and/or regions of interest are typically categorized into three main regimes corresponding to their characteristic initial relative velocity $v_{i}$:
  • On dwarf galaxy scales ($v_{i}\sim 10$ km/s), a cross-section per unit of DM mass in the range $\sigma/m_{\chi}\sim 0.1-10$ cm2/g enables efficient heat transfer within the core [74, 75, 21, 76, 77, 78]. This mechanism helps alleviate potential mass-deficit problems, i.e., the too-big-to-fail and core-vs-cusp problems (see e.g. [6]). While larger cross-sections can be considered, they eventually drive halos into gravothermal core collapse, yielding overly dense central regions [79]. Interestingly, this same process might account for the high central densities observed in certain compact or ultra-faint dwarf galaxies [80, 81, 82, 83].
  • On galaxy scales ($v_{i}\sim 100$ km/s), the interplay between the DM and baryons becomes non-negligible, rendering the constraints more subject to model dependencies. In Milky Way-like galaxies, very large cross-sections per unit of mass can lead to the evaporation of subhalos through frequent scatterings with host-halos DM particles. Such effects are expected to occur once $\sigma/m_{\chi}\gtrsim 5-10$ cm2/g [84, 81, 83]. For low surface brightness (LSB) galaxies, which are DM dominated, SIDM can account for the diversity in rotation curve shapes observed for $\sigma/m_{\chi}\gtrsim 3$ cm2/g [85, 86].
  • On galaxy cluster scales ($v_{i}\sim 1000$ km/s), the measurements of core densities from strong lensing gives the most strigent constraints [87], typically requiring cross-sections per unit of DM mass in the range $\sigma/m_{\chi}\lesssim 0.1-0.35$ cm2/g [88, 23]. Additional constraints arise from the analyses of spatial offsets between DM and galaxies in merging clusters—most notably the Bullet Cluster—where the observed alignment yields $\sigma/m_{\chi}\leq 1.25$ cm2/g [89]. Recent studies have further tightened these constraints to $\sigma/m_{\chi}\lesssim 0.2-0.5$ cm2/g [90, 91, 92].
@@
In the limit where the mediator mass is light compared to the transferred momentum, $m_{\phi,V}\ll\mu v_{i}$, (2.6) reduces to the Rutherford differential cross-section and presents a $1/v_{i}^{4}$ dependence. In the opposite limit where the interaction is carried by “light” mediators $\phi^{\prime}$ and $V^{\prime}$ that transmit a short-range force $m_{\phi^{\prime},V^{\prime}}\gg\mu v_{i}$, the corresponding cross-section scales as $1/m_{\phi^{\prime},V^{\prime}}^{4}$. Thus, in addition to introducing a velocity-dependent behavior at smaller mediator mass scales, SIDM mediated by a light boson leads to a desirable suppression of the cross-section in larger halo structures associated with larger characteristic values of $v_{i}$.
@@ 2.3 Validity of the Born regime
Throughout this work, we adopt a Born-level perturbative framework to describe the different dissipative processes. We are working in leading order in both, the DM self-interaction as well as in the emission of a single quantum. In this section, we specify the domain of applicability of our results.
We may write the dissipative $2\to 3$ differential cross section schematically as the product of the elastic scattering cross section and the probability of emission, $d\sigma_{2\to 3}\simeq\sigma_{2\to 2}dP_{1}$, a form that, for gauge emisson, finds itself well rooted in the soft-photon theorem (see Sec. 4). When $P_{1}\ll 1$, Bremsstrahlung can be seen as a perturbative correction to the $2\to 2$ DM elastic scattering process and the validity of the Born regime is determined by its validity in calculating elastic DM self-scattering. [^3] Exceptions are radiative bound state formation and, in the context of scalar $\phi^{3}$ theory, a strong infrared sensitivity leading to $P_{\rm em}=\mathcal{O}(1)$ already in the parameter region of interest; see App. B.2. In the NR framework associated with small-scale structure (see Sec. 2.2), the scalar and gauge mediating force between two DM candidates can be expressed as a Yukawa potential,
$V(r)=\frac{a_{1}a_{2}e^{-m_{\phi,V}r}}{4\pi r},$ (2.7)
where $a_{j}$ denotes the coupling of the mediator to DM candidates, i.e., $g_{j},y_{j}$ or $A_{j}/m_{j}$, with $\operatorname{sgn}{(a_{1}a_{2})}=-1\ (+1)$ being an attractive (repulsive) interaction. The Born regime is controlled by a small parameter whose form depends on whether the mediating force is short- or long-ranged.
@@
$b=\frac{\left|a_{1}a_{2}\right|\mu}{2\pi m_{\phi^{\prime},V^{\prime}}}\ll 1\quad(\text{Born regime}).$ (2.8)
Up to order unity factors, it may be seen as the ratio of the range of the potential $r_{\phi^{\prime},V^{\prime}}=1/m_{\phi^{\prime},V^{\prime}}$ to the would-be Coulombic Bohr radius $a_{0}=1/(\mu|a_{1}a_{2}|)$. The condition (2.8) imposes an lower bound on the heavy-mediator mass. [^4] Additional constraints arise from the NR framework we adopt, futher restricting the admissible range of $m_{\phi^{\prime},V^{\prime}}$; see Sec. 3.2 for details.
@@
Throughout this work we restrict our attention to real, single-emission tree-level bremsstrahlung from massive particles. In our case, soft enhancements arise solely from an energy integral over the single vector differential emission probability, $\int_{E}^{Q}dP_{1}/d\omega$, where $Q\lesssim\mu v_{i}^{2}/2$ denotes the hard scale of the process and $E$ is the IR scale. This integral equals the mean number $\lambda$ of unresolved vectors with energies $E\leq\omega\leq Q$, yielding an approximate Poisson multiplicity, $P_{n}=(\lambda^{n}/n!)e^{-\lambda}$. Using the eikonal factor $\mathcal{E}_{V}(q,\varepsilon)$ for emission with four-momentum $q$ (energy $\omega=q^{0}$ and three-momentum $\mathbf{q}$), polarization $\varepsilon$ from any external leg, Eq. (4.1), we obtain
@@
Here, $\mathcal{F}(\bm{\hat{}}{\mathbf{\mathbf{q}}})\equiv\sum_{\mathrm{pol}}\left|\omega\mathcal{E}_{V}(q,\varepsilon)\right|^{2}$ is the energy-independent, polarization-summed, angular-only kernel that integrates to $\omega^{2}$ times the dipole and the quadrupole emission factors, given in (4.8) and (4.13) respectively. The Born (single-emission) regime, in which virtual soft corrections can be neglected and bremsstrahlung is a perturbative correction to elastic $2\to 2$ process corresponds to $\lambda\ll 1$. For $\lambda\gtrsim 1$ multiple emissions are common and a resummed treatment is required. For scalar emission, the absence of gauge symmetry implies that there is no universal soft theorem. Nonetheless, the leading-order cancellation of soft divergences in real and virtual terms parallels the vector case [95]. In Sec. 4 and App. D we discuss the difference with vector emission at the level of the factorization of the amplitude.
@@
3 Energy differential emission cross-sections
In order to prepare for the expression of the energy differential cross-section and to set some of the notation, we begin this section by identifying the dissipative channels associated with each dSIDM model and their relevance in the kinematic regimes under consideration, i.e., NR colliding DM candidates with long-range and short-range mediation. We then introduce the terminology used throughout this work to distinguish between two distinct emission mechanisms: dipole and quadrupole radiation. With these definitions in place, we conclude this section by presenting the expressions for the differential energy cross-sections.

### v2

@@ Abstract
We derive the energy-differential cross section and energy loss rate for dissipative self-interacting dark matter (dSIDM) models within the Born regime using perturbative quantum field theory. Six dissipative scenarios are considered, incorporating the emission of particles that may be either massless or possess a kinematically allowed light mass. Both short-range and long-range force-mediated dSIDM interactions are examined. In the non-relativistic regime, we obtain closed-form expressions of the energy-differential cross sections by a controlled expansion in the initial relative dark matter velocity. Up to trivial factors, the leading-order squared emission amplitude is model-independent for massless emissions. Model dependence arises for massive particle emission and at the next-to-leading order. The latter reduces to three distinct cases. The derived analytical expressions exhibit excellent agreement with numerical computations, providing simple, ready-to-use formulas. Furthermore, we analyze the behavior of these processes in the soft emission limit. Our results show that additional corrections are necessary when applying factorization at the next-to-leading order in a velocity expansion to ensure consistency between the soft energy-differential cross section and the full counterparts across a broad energy range. Finally, we investigate the regime of perturbative validity in terms of the model parameters, identifying the conditions under which our results are applicable.
@@ 1 Introduction
Relaxing the assumption of a collisionless DM component also opens the door to additional phenomena, particularly dissipation, additionally modifying the thermodynamics of halos: radiative channels shorten the cooling time relative to dynamical and relaxation times, enabling contraction and catalyzing collapse to black holes [31, 32, 33, 34, 35, 36], compact substructure [37, 38, 39], or the formation of rotationally supported dark disks [40, 41, 42, 43]. Because of constraints on the self-interaction cross sections $\sigma/m$, typically only a fraction of the total DM is assumed to participate in such dissipative sector. Among other models, a well-studied dissipative realization is Atomic Dark Matter (ADM), which mimics the SM baryonic sector by allowing for the formation of hydrogen-like bound states [44, 45, 46, 47, 48, 49, 50]. Although ADM offers a rich phenomenological playground, much of its astrophysical impact stems from the net energy loss within halos, rather than from the specific details of the atomic spectrum. Consequently, some minimal dissipative models have been put forward in which the DM is simply composed of at most two DM candidates (fermionic or scalar) associated with either a massless, unbroken dark photon $U(1)_{D}$ [51, 52, 53], or through a massive dark gauge boson [54, 55]. These stripped-down setups already offer valuable insight into small-scale structure formation, avoiding the large parameter space inherent to more complex models. For additional works on dissipative SIDM (dSIDM) see [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66] and references therein.
@@
In the SM, the bremsstrahlung process corresponds to the emission of a photon in the scattering of two electrically charged particles, and is one of the most fundamental radiative processes. Two distinct cases arise. First, when the particles have different charge-to-mass ratios, dipole emission occurs. This scenario was first treated classically by Kramers [67], followed by quantum mechanical (Born) calculations by Oppenheimer [68] and Sugiura [69], and later by Sommerfeld who derived the exact non-relativistic solution using confluent hypergeometric functions [^1] To avoid the cumbersome exact solution, Elwert introduced a multiplicative correction that extends the Born approximation [70]. [71], a treatment now standard in advanced textbooks. More recently, dipole emissions have been revisited in [72]. Second, when the particles share the same mass-to-charge ratio, leading to the vanishing of the dipole piece and the dominance of quadrupole radiation. The corresponding cross section was first derived in a Born calculation by Gould [73, 74], and revisited recently in a calculation exact at all orders in the Coulomb interaction of two non-relativistic particles in [75]. Beyond the full energy spectrum, the soft-emission limit is also of particular interest: in this regime, for photon and graviton, the amplitude factorizes into a universal soft factor and a process-dependent hard amplitude [76].
@@
Bremsstrahlung-like processes are also found in many astrophysical and beyond the SM contexts. In stellar environments, similar types of radiation can arise from axion emission [77, 78], neutrino pair production [79, 80, 81], or scalar emission [82, 83]. Unsurprisingly, these multipole emissions also reappear in dissipative dark-sector settings: dipole bremsstrahlung governs cooling in ionized dark-electron–dark-proton plasmas, while quadrupole emission dominates in dark-electron–dark-electron interactions. However, a comprehensive theoretical framework for bremsstrahlung in generic dark sectors, departing from a simple gauge boson emission that includes dipole and higher multipoles emission, distinguishable and identical scatterers, different emitted quanta, and soft emission, has yet to be fully established.
@@
Motivated by this gap and the diversity of the SIDM proposals, we establish in this work a set of representative minimal dissipative scenarios dominated by bremsstrahlung emission. Our DM candidates include fermions (Dirac or Majorana) and scalars (real or complex), coupled through either scalar or vector mediator. Two distinct species are included in each scenario, allowing the study of both distinguishable and identical scatterers, and both the long-range (light mediator) and short-range (heavy mediator) regimes are covered. We work throughout in the non-relativistic (NR) limit, appropriate for small-scale structures where the characteristic DM relative velocity is $v_{i}\ll 1$. Within this framework, analytic formulae are derived for the energy-differential cross section and corresponding energy loss rate in every scenario. The results are compiled into a single catalogue and validated against numerical calculations and existing literature. Where previous studies relied on process-specific approximations, this catalogue offers a single lookup table covering the parameter space for the most relevant SIDM scenarios, and ready-to-use expressions that can be readily adapted to any combination of couplings, masses, and mediator properties. The present study is restricted to the perturbative Born regime; a detailed investigation of non-perturbative effects will be addressed in a future work.
@@
The paper is organized as follows: In Sec. 2, we introduce the set of dissipative self-interacting dark matter (dSIDM) scenarios we consider in this work, summarize the main existing constraints on the corresponding SIDM parameters, and discuss the domains of validity of our results. In Sec. 3, treating separately long-range and short-range mediation, we first identify the dissipative channels that appear in the Born (perturbative) regime and analyse their relevance in the NR limit, making explicit the emergence of a hierarchy of scales among them. Next, we provide a concise textbook review of NR radiation, distinguishing dipole from quadrupole emission. The section concludes with explicit expressions for the squared emission amplitudes and the energy-differential cross sections, both for distinguishable and identical particle scattering. In Sec. 4, we investigate the low-energy behavior of the different emission amplitudes and energy-differential cross sections obtained. We employ the standard factorisation that expresses the former as an elastic part multiplied by an emission piece, and discuss the limitations of such a method. We conclude this paper with Sec. 5, where we derive energy loss rates. The Appendices provide further details on our calculations, including the case where the emitted particle carries a small, kinematically allowed mass.
@@ 2.1 Field content and considered interactions
Again, we have a dimensionful trilinear interaction constant $A_{\phi}$ for the real scalar mediator and self-interaction quartic couplings $\lambda_{\phi}$ and $\lambda_{S}$. We take all couplings as real. [^2] We exclude off-diagonal couplings such as $\lambda_{12}S_{1}S_{2}\phi$, inducing a decay channel $S_{1}\to S_{2}\phi$ (or $S_{2}\to S_{1}\phi$ when $m_{1}>m_{2}+m_{\phi}$), altering the relative DM populations and introducing tangential complications.
A comment is in order about the Lagrangians (2.3), (2.4), and (2.5) when considering indistinguishable particle scattering, for which we then drop any generational index $j$ on the fields and couplings. In this case, we identify $S_{1}=S_{2}=S$ (and similarly for other fields) and it is understood that the sums over $j$ in (2.3) and (2.4) should be dropped. In other words, the normalization of coupling constants remains unchanged and does not receive a factor of two. Finally, the quartic scalar DM interaction term in (2.5) becomes $(1/4)\lambda_{S}|S|^{4}$ for $S$ and $(1/4!)\tilde{\lambda}_{S}|\tilde{S}|^{4}$ for $\tilde{S}$. It is for this reason that we have not included additional quartic $|S_{j}|^{4}$ interactions in (2.5).
@@
Of course, the possibilities considered above are not exhaustive, but still represent a fair fraction of the dSIDM possibilities. For simplicity, we shall also only consider exclusive cases where DM is of one principal fermionic or bosonic type. Mixed cases are, however, closely related since, as we shall see, the currents between fermions and scalars share many similarities in the NR limit. Notably absent from the above Lagrangians are axial-vector couplings and pseudoscalar couplings; in the latter case, the bremsstrahlung emission has been extensively studied in the context of axion emission from stars [77, 81, 78].
With the introduced field content, we identify six principal dSIDM combinations. For the processes where the interaction is mediated by a vector boson and involves the emission of a boson of the same type, two types of DM candidates are relevant for the corresponding dissipative processes: $\chi$ and $S$. These give rise to the dSIDM scenarios $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$, and $S_{1}S_{2}\to S_{1}S_{2}V$. In contrast, when the interaction is mediated by a scalar boson, and is associated with the emission of a scalar particle, the four DM candidates $S_{j}$, $\tilde{S}_{j}$, $\chi_{j}$, or $\tilde{\chi}_{j}$ are considered. These lead to the dSIDM scenarios $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$,  $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$ and $S_{1}S_{2}\to S_{1}S_{2}\phi$,  $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$.
@@ 2.2 Self-scattering constraints on model parameters
The success of SIDM lies in its ability to remain consistent across a wide range of astrophysical systems. In particular, the associated cross section per unit of DM mass ${\sigma}/{m_{\chi}}$ can be large enough to alter the inner density profiles of dwarf galaxies, yet small enough to avoid constraints on galaxy and galaxy cluster scales. Consequently, the principal bounds and/or regions of interest are typically categorized into three main regimes corresponding to their characteristic initial relative velocity $v_{i}$:
  • On dwarf galaxy scales ($v_{i}\sim 10$ km/s), a cross section per unit of DM mass in the range $\sigma/m_{\chi}\sim 0.1-10$ cm2/g enables efficient heat transfer within the core [85, 86, 21, 87, 88, 89]. This mechanism helps alleviate potential mass-deficit problems, i.e., the too-big-to-fail and core-vs-cusp problems (see e.g. [6]). While larger cross sections can be considered, they eventually drive halos into gravothermal core collapse, yielding overly dense central regions [90]. Interestingly, this same process might account for the high central densities observed in certain compact or ultra-faint dwarf galaxies [91, 92, 93, 94].
  • On galaxy scales ($v_{i}\sim 100$ km/s), the interplay between the DM and baryons becomes non-negligible, rendering the constraints more subject to model dependencies. In Milky Way-like galaxies, very large cross sections per unit of mass can lead to the evaporation of subhalos through frequent scatterings with host-halos DM particles. Such effects are expected to occur once $\sigma/m_{\chi}\gtrsim 5-10$ cm2/g [95, 92, 94]. For low surface brightness (LSB) galaxies, which are DM dominated, SIDM can account for the diversity in rotation curve shapes observed for $\sigma/m_{\chi}\gtrsim 3$ cm2/g [96, 97].
  • On galaxy cluster scales ($v_{i}\sim 1000$ km/s), the measurements of core densities from strong lensing gives the most stringent constraints [89], typically requiring cross sections per unit of DM mass in the range $\sigma/m_{\chi}\lesssim 0.1-0.35$ cm2/g [98, 23]. Additional constraints arise from the analyses of spatial offsets between DM and galaxies in merging clusters—most notably the Bullet Cluster—where the observed alignment yields $\sigma/m_{\chi}\leq 1.25$ cm2/g [20]. Recent studies have further tightened these constraints to $\sigma/m_{\chi}\lesssim 0.2-0.5$ cm2/g [22, 99, 100].
@@
In the limit where the mediator mass is light compared to the transferred momentum, $m_{\phi,V}\ll\mu v_{i}$, (2.6) reduces to the Rutherford differential cross section and presents a $1/v_{i}^{4}$ dependence. In the opposite limit where the interaction is carried by “light” mediators $\phi^{\prime}$ and $V^{\prime\mu}$ that transmit a short-range force $m_{\phi^{\prime},V^{\prime}}\gg\mu v_{i}$, the corresponding cross section scales as $1/m_{\phi^{\prime},V^{\prime}}^{4}$. Thus, in addition to introducing a velocity-dependent behavior at smaller mediator mass scales, SIDM mediated by a light boson leads to a desirable suppression of the cross section in larger halo structures associated with larger characteristic values of $v_{i}$.
@@ 2.3 Validity of the Born regime
Throughout this work, we adopt a Born-level perturbative framework to describe the different dissipative processes. We are working in leading order in both the DM self-interaction as well as in the emission of a single quantum. In this section, we specify the domain of applicability of our results.
We may write the dissipative $2\to 3$ differential cross section schematically as the product of the elastic scattering cross section and the probability of emission, $d\sigma_{2\to 3}\simeq\sigma_{2\to 2}dP_{1}$, a form that, for gauge emisson, finds itself well rooted in the soft-photon theorem (see Sec. 4). When $P_{1}\ll 1$, bremsstrahlung emission can be seen as a perturbative correction to the $2\to 2$ DM elastic scattering process, and the validity of the Born regime is determined by its validity in calculating elastic DM self-scattering. [^3] Exceptions are radiative bound state formation and, in the context of scalar $\phi^{3}$ theory, a strong infrared sensitivity leading to $P_{\rm em}=\mathcal{O}(1)$ already in the parameter region of interest; see App. B.2. In the NR framework associated with small-scale structure (see Sec. 2.2), the scalar and gauge mediating force between two DM candidates can be expressed as a Yukawa potential,
$V(r)=\zeta\frac{a_{1}a_{2}e^{-m_{\phi,V}r}}{4\pi r},$ (2.7)
where $a_{j}$ denotes the coupling of the mediator to DM candidates, i.e., $g_{j},y_{j}$ or $A_{j}/m_{j}$, with $\zeta=-1\ (+1)$ for a scalar (vector) mediator and $\zeta\operatorname{sgn}{(a_{1}a_{2})}=-1\ (+1)$ being an attractive (repulsive) interaction. The Born regime is controlled by a small parameter whose form depends on whether the mediating force is short- or long-ranged.
@@
$b=\frac{\left|a_{1}a_{2}\right|\mu}{4\pi m_{\phi^{\prime},V^{\prime}}}\ll 1\quad(\text{Born regime}).$ (2.8)
It may be seen parametrically as the ratio of the range of the potential $r_{\phi^{\prime},V^{\prime}}=1/m_{\phi^{\prime},V^{\prime}}$ to the would-be Coulombic Bohr radius $a_{0}=4\pi/(\mu|a_{1}a_{2}|)$. The condition (2.8) imposes a lower bound on the heavy-mediator mass. [^4] Additional constraints arise from the NR framework we adopt, further restricting the admissible range of $m_{\phi^{\prime},V^{\prime}}$; see Sec. 3.2 for details.
@@
Throughout this work, we restrict our attention to real, single-emission tree-level bremsstrahlung from massive particles. In our case, soft enhancements arise solely from an energy integral over the single vector differential emission probability, $\int_{E}^{Q}dP_{1}/d\omega$, where $Q\lesssim\mu v_{i}^{2}/2$ denotes the hard scale of the process and $E$ is the IR scale. This integral equals the mean number $\lambda$ of unresolved vectors with energies $E\leq\omega\leq Q$, yielding an approximate Poisson multiplicity, $P_{n}=(\lambda^{n}/n!)e^{-\lambda}$. Using the eikonal factor $\mathcal{E}_{V}(q,\varepsilon^{*})$ for emission with four-momentum $q$ (energy $\omega=q^{0}$ and three-momentum $\mathbf{q}$), polarization vector $\varepsilon^{\mu*}(q)$ from any external leg, Eq. (4.1), we obtain
@@
Here, $\mathcal{F}(\bm{\hat{}}{\mathbf{\mathbf{q}}})\equiv\sum_{\mathrm{pol}}\left|\omega\mathcal{E}_{V}(q,\varepsilon^{*})\right|^{2}$ is the energy-independent, polarization-summed, angular-only kernel that integrates to $\omega^{2}$ times the dipole and the quadrupole emission factors. The corresponding angular averages of these emission factors are given in (4.8) and (4.13), respectively. The Born (single-emission) regime, in which virtual soft corrections can be neglected, and bremsstrahlung is a perturbative correction to the elastic $2\to 2$ process corresponds to $\lambda\ll 1$. For $\lambda\gtrsim 1$ multiple emissions are common, and a resummed treatment is required. For scalar emission, the absence of gauge symmetry implies that there is no universal soft theorem. Nonetheless, the leading-order cancellation of soft divergences in real and virtual terms parallels the vector case [102]. In Sec. 4 and App. D we discuss the difference with vector emission at the level of the factorization of the amplitude.
@@
3 Energy-differential cross sections
In order to prepare for the expression of the energy-differential cross section and to set some of the notation, we begin this section by identifying the dissipative channels associated with each dSIDM model and their relevance in the kinematic regimes under consideration, i.e., NR colliding DM candidates with long-range and short-range mediation. We then introduce the terminology used throughout this work to distinguish between two distinct emission mechanisms: dipole and quadrupole radiation. With these definitions in place, we conclude this section by presenting the expressions for the differential energy cross sections.